# Interpretability of linear regression models of glassy dynamics

### v1

@@ Abstract
Data-driven models can accurately describe and predict the dynamical properties of glass-forming liquids from structural data. Accurate predictions, however, do not guarantee an understanding of the underlying physical phenomena and the key factors that control them. In this paper, we illustrate the merits and limitations of linear regression models of glassy dynamics built on high-dimensional structural descriptors. By analyzing data for a two-dimensional glass model, we show that several descriptors commonly used in glass‑transition studies display multicollinearity, which hinders the interpretability of linear models. Ridge regression suppresses some of the shortcomings of multicollinearity, but its solutions are not succinct enough to be physically interpretable. Only by using dimensional reduction techniques we eventually obtain linear models that strike a balance between prediction accuracy and interpretability. Our analysis points to a key role of local packing and composition fluctuations in the glass model under study.
@@ I Introduction
As the temperature of glass-forming liquids decreases, the relaxation timescale increases dramatically, while structural changes remain relatively modest [1, 2, 3]. Moreover, the dynamics exhibits pronounced spatial fluctuations, characterized by regions of high and low mobility, referred to as dynamic heterogeneities [4, 5, 6]. Despite the vivid patterns of dynamic heterogeneities, however, static snapshots of the system appear homogeneous and lack distinct features, at least to the naked eye. This has been confirmed through direct observations in computer simulations and colloidal glass experiments [7, 8]. To identify subtle but significant structural changes linked to dynamics, various structural order parameters based on physical intuition have been investigated [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Similar efforts have been done to identify the structural origins of plastic events in amorphous solids under loading [19, 20, 21, 22].
@@
Remarkably, recent studies have demonstrated that simple linear models, when combined with domain knowledge of glassy dynamics, such as coarse-graining techniques, can describe dynamic heterogeneities with remarkable accuracy [31, 46]. In some cases, their accuracy is comparable to that of more complex deep learning models. This outcome is particularly desirable, because it is often believed that simpler models not only provide accurate predictions at a minimal computational cost, but can also offer greater interpretability [47]. We caution, however, that employing a linear model per se is not sufficient to extract meaningful physical information: in high‑dimensional settings, linear models are often plagued by numerical instabilities, due to the so-called multicollinearity [48], that can obscure interpretation. Moreover, their solution may still be too high-dimensional to convey a physical meaning. The aim of this study is to address these issues in linear models for glassy dynamics and introduce remedies that restore interpretability.
To tackle the issue of interpretability in a rigorous manner, we will study quantitatively the consequences of multicollinearity in linear regression models of glassy dynamics, using a two-dimensional glass-forming liquid model. First, we will demonstrate how multicollinearity leads to instability in weight estimation within linear regression, thereby hindering the interpretation of feature importance. These problems affect several structural descriptors used in recent glass transition studies. Next, we will explore strategies to mitigate the effects of multicollinearity and to identify a set of low-dimensional linear models that achieve good performance accuracy with a minimum of structural information. We will critically evaluate and discuss the advantages and limitations of each approach, providing a comprehensive assessment of linear models of the structure-dynamics relationship in a glass-forming liquid model.
The paper is organized as follows. Section II defines the problem under investigation. Section III describes the physical model and our structure-dynamics dataset. Sections IV and Section V introduce simple linear regression models of glassy dynamics and demonstrate the effects of multicollinearity on the instability of the estimated weights. Section VI examines feature selection and extraction approaches that cope with multicollinearity while also reducing the dimensionality of the problem. Sections VII and VIII offer a critical outlook on our results and summarize the key findings of the work.
@@ II Problem statement
Typically, multicollinearity is not a significant concern in machine learning studies, as the primary objective is to improve the prediction accuracy. However, it becomes problematic when interpreting regression models in terms of feature importance. In fact, if the weight estimation is unstable, interpretation becomes unreliable. The weight estimate must be stable against fluctuations in the dataset or small changes in the hyperparameters. Such robustness is, of course, a basic requirement for interpretability of data-driven models, see Ref. 50 for a recent review on this topic. In our view, interpretable models must also provide a succinct description of the relationships between the variables of interest. This is in line with the physicists’ expectation that good models significantly compress the information involved in the problem at hand [51]. Phenomenological models in liquid state theory or statistical physics, for instance, are often based on a handful of independent variables – think of the hydrodynamic description of liquids [52], two-state models [53], etc. Hence, to be physically interpretable, data-driven models must be able to pinpoint a robust relationship between a small number of relevant variables, possibly carrying an intuitive physical meaning.
@@
To close this section, a comment on terminology is in order. To evaluate the quality of a data-driven model, one computes the estimate $\hat{\bf Y}$ for a set of input data that were not used to train the model. In the following, we will do that by splitting the original data set into a training and test set, as is customary. The estimate $\hat{\bf Y}$ is then considered as prediction for the test set; standard goodness-of-fit metrics can be used to evaluate their quality. From a conventional physicist’s viewpoint, this is a weak form of prediction: in the problem we will be dealing with, the data points will be sampled at precisely the same physical conditions both in the train and test set and no attempt will be made to extrapolate to different conditions, e.g., different temperatures or time scales. In this work, we will nonetheless stick to the word “prediction” for consistency with previous work [24] and with the widespread usage in the machine learning context. Should the readers feel uncomfortable with it, they can mentally replace “prediction” with “description” in the following.
@@ III.1 Simulation model
We use a three-component glass-forming liquid model composed of small (S), medium (M), and large (L) particles in two spatial dimensions with periodic boundary conditions [54]. The model is a variant of a well-studied binary mixture model [55, 56, 56, 57], to which we add particles of type M with an intermediate character between S and L [58]. The interaction between two particles is described by the Lennard-Jones potential
$v_{\alpha\beta}(r)=4\epsilon_{\alpha\beta}\left[\left(\frac{\sigma_{\alpha\beta}}{r}\right)^{12}-\left(\frac{\sigma_{\alpha\beta}}{r}\right)^{6}\right],$
where $\alpha,\beta=\mathrm{S},\mathrm{M},\mathrm{L}$. The potential is modified to ensure that it is twice continuously differentiable at the cutoff, following Ref. [56].
@@
We perform Monte Carlo (MC) simulations using translational displacements [59]. The MC move consists in picking a particle at random and displacing it by a vector drawn randomly within a square box of linear size $\delta_{\rm max}=0.12$. The move is accepted on the basis of the Metropolis acceptance rule, which ensures the detailed balance condition. Although MC simulations do not possess a physical timescale, time $t$ can be measured in units of MC sweeps, each comprising $N$ attempts to perform the MC move. In the regime of slow glassy dynamics of interest in this work, the Monte Carlo dynamics behaves similarly to other types of physical dynamics, e.g., Newtonian and Brownian dynamics [60]. Therefore, we analyze the MC dynamics by following particle trajectories and calculating time-dependent observables as usual. The glassy dynamics of the model has been studied in Ref. [54] by computing the self intermediate scattering function. In this study, we focus on $T=0.30$, which is the lowest temperature at which we can equilibrate the system within our computational timescale.
@@ III.2 Dynamic propensity
To investigate the heterogeneity of glassy dynamics in real space and assess its connection with the static structure, we use the iso-configurational ensemble [49, 11]. A set of $n=10$ statistically uncorrelated configurations are obtained at equilibrium conditions at temperature $T=0.30$. From each of these equilibrium configurations, we generate an ensemble of trajectories using MC dynamics at $T=0.30$ using 30 different initial random seeds. We then compute the dynamic propensity
$p_{i}(t)=\left\langle|\Delta{\bf r}^{\rm CR}_{i}(t)|\right\rangle_{\rm iso},$
@@ III.3 Behler-Parrinello descriptor
To characterize the local structure around each particle, we use the Behler-Parrinello (BP) descriptor [62], which has been widely used to study structure-property relationships, including the description of glassy dynamics in two dimensions [23, 63]. The descriptor comprises two subsets of features that characterize radial and angular correlations, respectively. Following Refs. [23, 63], we further distinguish features according to the species of the neighboring particles.
@@
where $\theta_{ijk}$ is the angle at the corner $i$ of the triangle defined by particles $i$, $j$, and $k$, and $\xi$, $\lambda$, and $\zeta$ are parameters that are varied systematically. For each pair of species, $(\alpha,\beta)$, we employ the same set of $22$ parameters ($\xi$, $\lambda$, $\zeta$) given in Ref. [23] in unit of $\sigma_{\rm LS}$. The features $\Psi^{\alpha\beta}(k)$ are parametrized by an integer $k$ and the parameters $(\xi=\xi_{k},\lambda=\lambda_{k},\zeta=\zeta_{k})$ with $0\leq k\leq 21$ are shown in Table 1. Thus, we have $132(=6\times 22)$ angular features.
@@ III.4 Physically motivated descriptors
$u_{i}=\frac{1}{2}\sum_{j\neq i}v_{\alpha_{i}\beta_{j}}(r_{ij}),$
@@
$\Psi_{6,i}=\frac{1}{z_{i}}\left|\sum_{j=1}^{z_{i}}e^{\sqrt{-1}\ 6\theta_{ij}}\right|,$
where $\theta_{ij}$ is the angle between ${\bf r}_{ij}={\bf r}_{j}-{\bf r}_{i}$ and the x-axis. The nearest neighbors are again defined as those within $r_{ij}<1.5\sigma_{\alpha_{i}\beta_{j}}$. $\Psi_{6,i}$ quantifies hexagonal order, taking the value $1$ for perfect hexagonal packings and smaller values for disordered packings [65, 66].
@@
$\Theta_{i}=\frac{1}{z_{i}}\sum_{\langle jk\rangle}\left|\theta_{jk}-\theta_{jk}^{\rm ref}\right|,$
@@
$\overline{\varphi}_{i}(\ell)=\sum_{j\in\mathcal{N}_{i}}(\sigma_{\alpha_{i}\beta_{j}})^{2}e^{-r_{ij}/\ell}.$
@@
We also use the JBB descriptor introduced in the paper by Jung, Biroli, and Berthier [26]. This descriptor is based on $\overline{\rho}_{i}$, $u_{i}$, and $\pi_{i}$. In this work, we ignore the variance of the potential energy, which was included in Ref. [26]. The JBB descriptor incorporates particle‑species information more explicitly, as shown below. A first set of coarse-grained features is obtained using the same procedure as for the SLO descriptor, considering the whole set of neighbors, irrespective of their species, with 10 different coarse-graining length scales $\ell$. In addition, the JBB descriptor includes coarse-grained features obtained by a procedure similar to Eq. (6) but taking species into account:
@@
$\overline{\rho}_{i}^{\alpha}(\ell)=\sum_{j\in\mathcal{N}_{i}^{\alpha}}e^{-r_{ij}/\ell},$
and the sums are restricted to neighbors of species $\alpha$. Thus, in addition to the species-independent coarse-graining defined by Eq. (6), we consider coarse-graining based on Eq. (7) using the three types of particles ($\alpha=$ S, M, and L). This yields a total of $M=120$ features (3 descriptors $\times$ 4 types $\times$ 10 length scales).
@@ III.5 Dataset
$Y_{i}=\frac{p_{i}-{\mathbb{E}}[p]}{\sqrt{\mathrm{Var}[p]}},$
@@ III.6 Pearson correlation coefficients
We first compute the Pearson coefficients $R[{X^{(f)}},{Y}]$ between the dynamic propensity Y and each structural feature $\textbf{X}^{(f)}$ of the BP descriptor. Because of the very small dependence of the dynamic propensity on the species of the particles, we include all the particles, irrespective of their species, in the analysis of correlations. The Pearson coefficients are conveniently assembled in vector form
@@
In Fig. 1, we show results for ${\bm{R}}[{X},{Y}]$ obtained using the BP descriptor. The ordering of the features follows the logic described in Sec. III.3. We first separate the features into a radial and an angular sector. Then, within each of these two sectors, the features are grouped into blocks according to the species or species pair for which the feature is computed. We see that the correlations are generally modest ($|R|<0.5$) and spread over the whole set of structural features. In particular, both radial and angular features can have similar correlations or anticorrelations with the dynamic propensity. The presence of minima and maxima in each block of the radial features $G^{\alpha}$ can be connected to the peaks of the partial radial distribution functions, although we could not identify a straightforward interpretation. The angular sector displays systematic effects due to the chemical composition of the particle environment. Namely, features involving SS and LL pairs of neighbors are more strongly correlated or anti-correlated to the propensity. By contrast, $R$ shows no clear trend within each of the blocks of the angular sector.
The results for the two physically motivated descriptors introduced in Sec. III.4 display qualitatively similar trends, see the Appendix A. Among all the individual structural features entering our datasets, the steric order parameter $\overline{\Theta}_{i}(\ell)$, coarse-grained at $\ell=2.0$, exhibits the strongest correlation ($R\approx 0.65$) with the dynamic propensity, making it the best-performing single feature within this simple correlation analysis. While the local volume fraction $\overline{\varphi}_{i}(\ell)$ and the coordination number $\overline{z}_{i}(\ell)$ are negatively correlated with the dynamic propensity, as expected, the local number density $\overline{\rho}_{i}(\ell)$ is positively correlated to it. This counter-intuitive result is likely a non-trivial effect of composition fluctuations in our ternary glass model.
@@
In Fig. 2, we show the correlation matrix for the BP descriptor. It has a distinct block structure, displaying strong positive or negative correlations within subsets of features. Correlations within species-wise blocks are particularly pronounced for the angular features, see the bottom right region of Fig. 2. Moreover, there are non-trivial cross-correlations also between radial and angular features. A similar block structure is also apparent in the physically motivated descriptors, see the Appendix A. In those cases, however, block correlations are due to coarse-graining a given structural feature over a range of similar distances $\ell$. As we will demonstrate in Sec. IV, the redundancy of these structural descriptors has a significant negative impact on the linear modeling of the dynamic propensity.

### v2

@@ Abstract
Data-driven models can accurately describe and predict the dynamical properties of glass-forming liquids from structural data. Accurate predictions, however, do not guarantee an understanding of the underlying physical phenomena and the key factors that control them. In this paper, we illustrate the merits and limitations of linear regression models of glassy dynamics built on high-dimensional structural descriptors. By analyzing data for a two-dimensional glass model, we show that several descriptors commonly used in glass‑transition studies display multicollinearity, which hinders the interpretability of linear models. Ridge regression suppresses some of the shortcomings of multicollinearity, but its solutions are not concise enough to be physically interpretable. Only by using dimensional reduction techniques we do eventually obtain linear models that strike a balance between prediction accuracy and interpretability. Our analysis points to a key role of local packing and composition fluctuations in the glass model under study.
@@ I Introduction
As the temperature of glass-forming liquids decreases, the relaxation timescale increases dramatically, while structural changes remain relatively modest [1, 2, 3]. Moreover, the dynamics exhibits pronounced spatial fluctuations, characterized by regions of high and low mobility, referred to as dynamic heterogeneities [4, 5, 6]. Despite the vivid patterns of dynamic heterogeneities, however, static snapshots of the system appear homogeneous and lack distinct features, at least to the naked eye. This has been confirmed through direct observations in computer simulations and colloidal glass experiments [7, 8]. To identify subtle but significant structural changes linked to dynamics, various structural order parameters based on physical intuition have been investigated [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Similar efforts have been made to identify the structural origins of plastic events in amorphous solids under loading [19, 20, 21, 22].
@@
Remarkably, recent studies have demonstrated that simple linear models, when combined with domain knowledge of glassy dynamics, such as coarse-graining techniques, can describe dynamic heterogeneities with remarkable accuracy [31, 46]. In some cases, their accuracy is comparable to that of more complex deep learning models. This outcome is particularly desirable, because it is often believed that simpler models not only provide accurate predictions at a minimal computational cost, but can also offer greater interpretability [47]. We caution, however, that employing a linear model per se is not sufficient to extract meaningful physical information: in high‑dimensional settings, linear models are often plagued by numerical instabilities, because of so-called multicollinearity [48], that can obscure interpretation. Moreover, their solution may still be too high-dimensional to convey a physical meaning. The aim of this study is to address these issues in linear models for glassy dynamics and introduce remedies that restore interpretability.
To tackle the issue of interpretability in a rigorous manner, we will quantitatively study the consequences of multicollinearity in linear regression models of glassy dynamics, using a two-dimensional glass-forming liquid model. First, we will demonstrate how multicollinearity leads to instability in weight estimation within linear regression, thereby hindering the interpretation of feature importance. These problems affect several structural descriptors used in recent glass transition studies. Next, we will explore strategies to mitigate the effects of multicollinearity and to identify a set of low-dimensional linear models that achieve good performance accuracy with a minimum of structural information. We will critically evaluate and discuss the advantages and limitations of each approach, providing a comprehensive assessment of linear models of the structure-dynamics relationship in a glass-forming liquid model.
The paper is organized as follows. Section II defines the problem under investigation. Section III describes the physical model and our structure-dynamics dataset. Sections IV and V introduce simple linear regression models of glassy dynamics and demonstrate the effects of multicollinearity on the instability of the estimated weights. Section VI examines feature selection and extraction approaches that cope with multicollinearity while also reducing the dimensionality of the problem. Sections VII and VIII offer a critical outlook on our results and summarize the key findings of the work.
@@ II Problem statement
Typically, multicollinearity is not a major concern in machine-learning studies, where the primary objective is to optimize predictive accuracy. It becomes problematic, however, when regression models are interpreted in terms of feature importance. In particular, if the estimation of the regression weights is unstable, any physical interpretation becomes unreliable. Two distinct mechanisms can undermine the stability of the estimated weights because of multicollinearity. The first one corresponds to large dataset-to-dataset fluctuations: the values of the weights may vary significantly depending on the particular subsample of data used for training, even when all datasets are generated under identical physical conditions. In the large-sample regime, this variance is suppressed, leading to robust and reproducible weight estimates. A second, qualitatively different pathological behavior persists even in the large-sample limit. In this case, the weights exhibit oscillatory behavior across features. This instability originates also from multicollinearity. Note that this effect is not associated with statistical uncertainty but with the geometry of the feature space itself, and therefore cannot be mitigated by increasing the dataset size. Robustness of the estimated weights against both sources of instability is a basic requirement for interpretability of data-driven models.
The very notion of interpretability is context dependent [50] and remains the subject of active debate; see Ref. 51 for a recent review in the physics context. Recent works have emphasized the role of sparsity [52], whereby the relevant features of a data-driven model stand out clearly from the bulk, e.g., in terms of their weights or mutual distances in feature space. In our view, physically interpretable models must additionally provide a concise and robust description of the relationships between the variables of interest. This perspective is consistent with the physicists’ expectation that successful models strongly compress the information contained in the problem [53]. Phenomenological models in liquid-state theory and statistical physics, for instance, are often built from a small number of independent variables, such as in hydrodynamic descriptions or two-state models. Accordingly, to be physically interpretable, data-driven models must identify robust relationships involving a limited set of relevant variables, each carrying a clear physical meaning.
@@
To close this section, a comment on terminology is in order. To evaluate the quality of a data-driven model, one computes the estimate $\hat{\bf Y}$ for a set of input data that were not used to train the model. In the following, we will do that by splitting the original data set into a training and test set, as is customary. The estimate $\hat{\bf Y}$ is then considered as prediction for the test set; standard goodness-of-fit metrics can be used to evaluate their quality. From a conventional physicist’s viewpoint, this is a weak form of prediction: in the problem we will be dealing with, the data points will be sampled at precisely the same physical conditions both in the train and test set and no attempt will be made to extrapolate to different conditions. We will nonetheless stick to the word “prediction” throughout this work, for consistency with previous work [24] and with the widespread usage in the machine learning context. Should the readers feel uncomfortable with it, they can mentally replace “prediction” with “description” in the following. Only in Sec. VII we will assess the cross-state predictions of the models, by extrapolating the dynamic propensity to different temperatures and time scales.
@@ III.1 Simulation model
We use a three-component glass-forming liquid model composed of small (S), medium (M), and large (L) particles in two spatial dimensions with periodic boundary conditions [54]. The model is a variant of a well-studied binary mixture model [55, 56, 56, 57], to which we add particles of species M with an intermediate character between S and L [58]. The interaction between two particles is described by the Lennard-Jones potential
$v_{\alpha\beta}(r)=4\epsilon_{\alpha\beta}\left[\left(\frac{\sigma_{\alpha\beta}}{r}\right)^{12}-\left(\frac{\sigma_{\alpha\beta}}{r}\right)^{6}\right],$ (2)
where $\alpha,\beta=\mathrm{S},\mathrm{M},\mathrm{L}$. The potential is modified to ensure that it is twice continuously differentiable at the cutoff, following Ref. 56.
@@
We perform Monte Carlo (MC) simulations using translational displacements [59]. The MC move consists in picking a particle at random and displacing it by a vector drawn randomly within a square box of linear size $\delta_{\rm max}=0.12$. The move is accepted on the basis of the Metropolis acceptance rule, which ensures the detailed balance condition. Although MC simulations do not possess a physical timescale, time $t$ can be measured in units of MC sweeps, each comprising $N$ attempts to perform the MC move. In the regime of slow glassy dynamics of interest in this work, the Monte Carlo dynamics behaves similarly to other types of physical dynamics, e.g., Newtonian and Brownian dynamics [60]. Therefore, we analyze the MC dynamics by following particle trajectories and calculating time-dependent observables as usual. The glassy dynamics of the model has been studied in Ref. 54 by computing the self intermediate scattering function. In this study, we mainly focus on $T=0.30$, which is the lowest temperature at which we can equilibrate the system within our computational timescale.
@@ III.2 Dynamic propensity
To investigate the heterogeneity of glassy dynamics in real space and assess its connection with the static structure, we use the isoconfigurational ensemble [49, 11]. A set of $n=10$ statistically uncorrelated configurations are obtained at equilibrium conditions at temperature $T=0.30$. From each of these equilibrium configurations, we generate an ensemble of trajectories using MC dynamics at $T=0.30$ using 60 different initial random seeds. We then compute the dynamic propensity
$p_{i}(t)=\left\langle|\Delta{\bf r}^{\rm CR}_{i}(t)|\right\rangle_{\rm iso},$ (3)
@@ III.3 Behler-Parrinello descriptor
To characterize the local structure around each particle, we use the Behler-Parrinello (BP) descriptor [62], which has been widely used to study structure-property relationships, including the description of glassy dynamics in two dimensions [23, 63]. The descriptor comprises two subsets of features that characterize radial and angular correlations, respectively. Following Refs. 23, 63, we further distinguish features according to the species of the neighboring particles.
@@
where $\theta_{ijk}$ is the angle at the corner $i$ of the triangle defined by particles $i$, $j$, and $k$, and $\xi$, $\lambda$, and $\zeta$ are parameters that are varied systematically. For each pair of species, $(\alpha,\beta)$, we employ the same set of $22$ parameters ($\xi$, $\lambda$, $\zeta$) given in Ref. 23 in unit of $\sigma_{\rm LS}$. The features $\Psi^{\alpha\beta}(k)$ are parametrized by an integer $k$ and the parameters $(\xi=\xi_{k},\lambda=\lambda_{k},\zeta=\zeta_{k})$ with $0\leq k\leq 21$ are shown in Table 1. Thus, we have $132(=6\times 22)$ angular features.
@@ III.4 Physically motivated descriptors
$u_{i}=\frac{1}{2}\sum_{j\neq i}v_{\alpha_{i}\beta_{j}}(r_{ij}),$ (8)
@@
$\Psi_{6,i}=\frac{1}{z_{i}}\left|\sum_{j=1}^{z_{i}}e^{\sqrt{-1}\ 6\theta_{ij}}\right|,$ (9)
where $\theta_{ij}$ is the angle between ${\bf r}_{ij}={\bf r}_{j}-{\bf r}_{i}$ and the $x$ axis. The nearest neighbors are again defined as those within $r_{ij}<1.5\sigma_{\alpha_{i}\beta_{j}}$. $\Psi_{6,i}$ quantifies hexagonal order, taking the value $1$ for perfect hexagonal packings and smaller values for disordered packings [65, 66].
@@
$\Theta_{i}=\frac{1}{z_{i}}\sum_{\langle jk\rangle}\left|\theta_{jk}-\theta_{jk}^{\rm ref}\right|,$ (10)
@@
$\overline{\varphi}_{i}(\ell)=\frac{1}{\overline{\rho}_{i}(\ell)}\sum_{j\in\mathcal{N}_{i}}(\sigma_{\alpha_{i}\beta_{j}})^{2}e^{-r_{ij}/\ell}.$ (12)
@@
We also use the JBB descriptor introduced in the paper by Jung, Biroli, and Berthier [26]. This descriptor is based on $\overline{\rho}_{i}$, $u_{i}$, and $\pi_{i}$. In this work, we ignore the variance of the potential energy, which was included in Ref. 26. The JBB descriptor incorporates particle‑species information more explicitly, as shown below. A first set of coarse-grained features is obtained using the same procedure as for the SLO descriptor, considering the whole set of neighbors, irrespective of their species, with 10 different coarse-graining length scales $\ell$. In addition, the JBB descriptor includes coarse-grained features obtained by a procedure similar to Eq. (13) but taking species into account:
@@
$\overline{\rho}_{i}^{\alpha}(\ell)=\sum_{j\in\mathcal{N}_{i}^{\alpha}}e^{-r_{ij}/\ell},$ (15)
and the sums are restricted to neighbors of species $\alpha$. Thus, in addition to the species-independent coarse-graining defined by Eq. (13), we consider coarse-graining based on Eq. (14) using the three species of particles ($\alpha=$ S, M, and L). This yields a total of $M=120$ features (3 descriptors $\times$ 4 types $\times$ 10 length scales).
@@ III.5 Dataset
$Y_{i}=\frac{p_{i}-{\mathbb{E}}[p]}{\sqrt{\mathrm{Var}[p]}},$ (16)
@@ III.6 Pearson correlation coefficients
We first compute the Pearson coefficients $R[{X^{(f)}},{Y}]$ between the dynamic propensity Y and each structural feature $\textbf{X}^{(f)}$ of the BP descriptor. Because of the weak dependence of the dynamic propensity on the species of the particles, see the Appendix A, we include all the particles, irrespective of their species, in the analysis of correlations. The Pearson coefficients are conveniently assembled in vector form
@@
Note that, because of the finite number of isoconfigurational trajectories, there is an upper bound on the Pearson coefficient between the propensity Y and any structural feature [27]. We used the method described in Ref. 27 to compute the maximum Pearson coefficient $R_{\textrm{max}}$ that can be achieved with the current statistics. We found $R_{\textrm{max}}\approx 0.98$, which is fully satisfactory for our purposes.
In Fig. 1, we show results for ${\bm{R}}[{X},{Y}]$ obtained using the BP descriptor. The ordering of the features follows the logic described in Sec. III.3. We first separate the features into a radial and an angular sector. Then, within each of these two sectors, the features are grouped into blocks according to the species or species pair for which the feature is computed. We see that the correlations are generally modest ($|R|<0.5$) and spread over the whole set of structural features. In particular, both radial and angular features can have similar correlations or anticorrelations with the dynamic propensity. The presence of minima and maxima in each block of the radial features $G^{\alpha}$ can be connected to the peaks of the partial radial distribution functions, although we could not identify a straightforward interpretation. The angular sector displays systematic effects because of the chemical composition of the particle environment. Namely, features involving SS and LL pairs of neighbors are more strongly correlated or anti-correlated to the propensity. By contrast, $R$ shows no clear trend within each of the blocks of the angular sector.
The results for the two physically motivated descriptors introduced in Sec. III.4 display qualitatively similar trends, see the Appendix B. Among all the individual structural features entering our datasets, the steric order parameter $\overline{\Theta}_{i}(\ell)$, coarse-grained at $\ell=2.0$, exhibits the strongest correlation ($R\approx 0.65$) with the dynamic propensity, making it the best-performing single feature within this simple correlation analysis. While the local volume fraction $\overline{\varphi}_{i}(\ell)$ and the coordination number $\overline{z}_{i}(\ell)$ are negatively correlated with the dynamic propensity, as expected, the local number density $\overline{\rho}_{i}(\ell)$ is positively correlated to it. This counter-intuitive result is likely a nontrivial effect of composition fluctuations in our ternary glass model.
@@
In Fig. 2, we show the correlation matrix for the BP descriptor. It has a distinct block structure, displaying strong positive or negative correlations within subsets of features. Correlations within species-wise blocks are particularly pronounced for the angular features, see the bottom right region of Fig. 2. Moreover, there are non-trivial cross-correlations also between radial and angular features. A similar block structure is also apparent in the physically motivated descriptors, see the Appendix B. In those cases, however, block correlations are due to coarse-graining a given structural feature over a range of similar distances $\ell$. As we will demonstrate in Sec. IV, the redundancy of these structural descriptors has a significant negative impact on the linear modeling of the dynamic propensity.